OpenStudy (anonymous):

Would you expect to observe the Compton effect more readily with scattering targets composed of atom with high atomic number or those composed of atoms with low atomic number? Explain!

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

@jhonyy9

OpenStudy (calculusfunctions):

@gerryliyana , I see you asked for my help four days ago, but I'm just seeing your question now. I hope it's still helpful. In short, when a photon collides with a particle, an electron for example, at an angle θ, the scattered photon will now possess less energy and a lower frequency, as a consequence. Also it changes momentum. Since it experiences a decreases in frequency, it's wavelength increases. If a photon with a given wavelength collides with a free electron at rest, at an angle θ, then it's scattered wavelength increases, and\[\lambda _{2}-\lambda _{1}=\frac{ h }{ m _{e}c }(1-\cos \theta )\]where the Δλ is the change in wavelength, m sub e is the mass of the free electron at rest, and h = 6.626 x 10 ˆ-34 is Planck's constant. From the equation you can observe that as the mass of the electron increases, the wavelength decreases.